# Automatic Fused Multimodal Deep Learning for Plant Identification

### v1

@@ 3.2 Finding Final Architecture
layers fused from flower, leaf, fruit and stem unimodal models respectively. $A$ – first activation,
$O$ – global average pooling on the top of a model, $RN$$N$-th inverted residual block.
@@
F1macro values have been achieved by models for 2 epochs during the search (Section 2.4.2).
@@
Figure 4: Validation Accuracies of Sampled Architectures per Epoch
@@ 3.3 Final Model
To train the final model, we adjust various hyperparameters of the first architecture from Table 2. The number of neurons per fusion layer is increased to 512, and to enhance generalization, dropout rates of 0.1, 0.2, 0.2, and 0.1 are set for the fusion layers and the classification layer, respectively. Additionally, we apply L2 regularization of $10^{-5}$, weight decay of 0.01, and label smoothing of 0.02. Batch normalization was added after each fusion layer.
With all the aforementioned changes, the model is set up for training for 100 epochs, with Adam optimizer, initial learning rate of 0.001 with exponential decay of 0.95 per epoch, and early stopping with patience of 10, resulting in a validation accuracy of 81.72%, as shown in Figure 5. To further enhance performance, we fine-tune the resulting model by unfreezing unimodal base models. To do so, all hyperparameters are retained, except for the initial learning rate, which has been replaced with a value of $10^{-4}$ to prevent erasing previously learned knowledge. As a result, as shown in Figure 5, the validation accuracy increases to 83.45%.
Figure 5: Validation Accuracy of Final Model per Epoch
Blue – before fine-tuning, green – after fine-tuning.
@@ 3.4 Comparison with the Baseline
We collect the performance metrics of the proposed model and the baseline on the test set, which are displayed in Table 3.
Proposed Model Baseline
Accuracy 0.8348 0.7241
Top-5 accuracy 0.9418 0.8891
Top-10 accuracy 0.9604 0.9293
Precisionmicro 0.8348 0.7241
Precisionmacro 0.8074 0.7220
Recallmicro 0.8348 0.7241
Recallmacro 0.7739 0.6408
F1micro 0.8348 0.7241
F1macro 0.7751 0.6529
Table 3: Performance Metrics of Proposed Model and Baseline
These results indicate that the proposed model significantly outperforms the baseline across all explored metrics and suggest that our automated fused model is substantially more effective and reliable in plant classification on the given dataset compared to a straightforward late fusion approach. The improvement is particularly notable in the macro recall, where it reaches +0.1331. This suggests that the proposed model misses the correct instances of a class significantly less frequently than the baseline. Conversely, a more modest improvement is observed in macro precision, where it reaches +0.0854. Interestingly, the difference between micro and macro precision implies that the baseline demonstrates relatively high precision when positively classifying instances of rare classes.
@@ 3.4.1 McNemar’s Test Results
We also conduct McNemar’s test to compare the performance of two multimodal plant classifiers, the proposed model and the baseline. The contingency table of model predictions is provided in Table 4. The test yielded a statistic $\chi^{2}_{1}=617.53$, which is statistically significant ($p<0.001$), indicating a significant difference in performance metrics between our proposed model and the baseline.
$n_{00}$: 1161 $n_{01}$: 248
$n_{10}$: 1192 $n_{11}$: 5927
@@
Table 4: Contingency Table of Proposed Model and Baseline
3.5 Comparison with Unimodal Models
@@
One can observe from Table 1 and Table 3 that both our proposed model and late fusion significantly outperform each individual unimodal model. However, these observations do not fully consider the fact that the unimodal models are evaluated on isolated unimodal datasets, whereas the multimodal model is evaluated on a multimodal dataset containing missing modalities and duplications.
To address this and in order to have a fair comparison, we evaluate both our proposed model and the unimodal models on isolated modalities from the multimodal dataset. The results of this comparison are presented in Table 5.
Unimodal Model Proposed Model
Flower 0.5638 0.4647
Leaf 0.3755 0.2094
Fruit 0.6158 0.4166
Stem 0.3622 0.2279
Table 5: Comparison of F1macro of Proposed and Unimodal Models on Individual Modalities
The metrics presented in Table 5 account for the absence of target modalities by including only those dataset instances that contain the required modality. While the multimodal model demonstrates commendable performance across individual modalities, it generally underperforms compared to the unimodal models when restricted to using only one modality. In detail, interestingly, the performance gap between the multimodal and unimodal models is notably smaller for the flower modality. This might be influenced by the abundance of flower examples in the dataset, providing a richer training environment. Additionally, factors such as the number of classes within each modality and the inherent discriminative capacity of the organs likely affect the performance metrics of the multimodal model. The observed lower performances could stem from a deficiency in data, whereas discrepancies with the unimodal models might be attributed to interference from other modal components within the multimodal system. Given these observations, it is expected that models specifically trained on individual modalities would surpass the performance of a multimodal model when only a single modality is accessible.
3.6 Robustness to Missing Modalities
While cases where a single or all organs are available, represent two extremes, there exists a considerable number of image combinations within the dataset where 2–3 modalities are available. Therefore, it is essential to evaluate the performance of our proposed model and the baseline on all possible subsets of 2–3 modalities. We present the results of this evaluation in Table 6. Similarly to Table 5, it is important to note that the metrics in Table 6 account for the absence of target modalities by excluding instances that miss any of the required modalities.
Modalities Proposed Model Baseline
Flower, Leaf 0.6236 0.5102
Flower, Fruit 0.6233 0.5907
Flower, Stem 0.4562 0.4361
Leaf, Fruit 0.6097 0.5818
Leaf, Stem 0.5356 0.4603
Fruit, Stem 0.6142 0.5382
Flower, Fruit, Leaf 0.7816 0.5985
Flower, Stem, Leaf 0.7332 0.5261
Flower, Fruit, Stem 0.6079 0.5416
Leaf, Fruit, Stem 0.7638 0.6143
@@ 3.6 Robustness to Missing Modalities
Table 6: Comparison of F1macro of Proposed and Baseline Models on Subsets of Modalities
According to Table 5: (i) Across all cases, the proposed model consistently outperforms the baseline, indicating that our approach is superior to naive averaging in any multimodal configuration. (ii) In the cases of leaf–stem and fruit–stem, the latter exhibits higher performance in both models. This suggests that the discriminative features of fruits outweigh the representativeness of leaves. This observation aligns with the metrics presented in Table 1. (iii) An interesting observation arises from the comparison between the combinations of flower–leaf versus flower–fruit. It is notable that the former combination exhibits slightly better performance over the latter in the proposed model, whereas the opposite trend is observed in the baseline. This empirical evidence supports the notion put forth by Nhan et al. (2020) that flowers are the most discriminative organ. However, this discriminativeness seems to be insufficient when all modalities contribute equally to the final prediction, as in late fusion. In such cases, late fusion may favor another highly distinctive organ, such as fruits. In contrast, our model, employing a sophisticated, learned prediction strategy, demonstrates more robustness in either situation. It appears to prefer leaves, as they are more numerous, and the model makes fewer biased assumptions about the equality of modalities. Consequently, it more effectively analyzes combinations of modalities, leading to improved performance across various configurations. (iv) Additionally, the results suggest that stems may have the least positive impact on performance. This finding aligns with the observations in (Nhan et al., 2020), which argue that stems are the least discriminative organ. However, this particular conclusion should be interpreted with caution, as stems are also the least available organ in the dataset.
Overall, the proposed model demonstrates robustness to missing modalities. However, the availability of data remains a bottleneck for performance. Evidently, the less discriminative organs are, the more training data is needed to achieve higher results. This underscores the importance of data availability and diversity in training robust multimodal models.
@@ 4 Discussion
Our proposed multimodal DL model, leveraging plant organ images fused via the MFAS algorithm, proves highly effective in automating plant classification, reaching an accuracy of 83.48%. Its performance significantly surpasses the defined baseline and outperforms state-of-the-art models in various aspects with comparable settings, as summarized in Table 7. To be more specific, for example, Ghazi et al. (2017) employed all organs from the PlantCLEF2015 dataset. Despite utilizing larger and more sophisticated classifiers, along with advanced scoring methods, they achieved an accuracy of 80.18%. Despite our architecture being relatively small (6 million parameters), the incorporation of multimodal fusion allowed us to outperform their model, highlighting the strength of multimodality and the advantages of an optimized fusion strategy of our proposed method. Similarly, de Lutio et al. (2021) sampled a large dataset comprising 56608 images of 977 species from the iNaturalist database. Utilizing images and spatio-temporal context, their multimodal model achieved accuracies of 79.12% without satellite imagery and 79.73% with it. While a direct comparison is challenging due to the use of different datasets, our approach, incorporating various organ modalities, yielded higher results. Furthermore, Lee et al. (2018), also using the PlantCLEF2015 dataset, achieved an accuracy of 68.5%. Ge et al. (2016) and Nguyen et al. (2016) extracted flower species from the PlantCLEF2015 dataset and achieved an accuracy of 52.1% and 67.45% respectively. All these results are significantly lower than ours, underscoring the effectiveness of our approach.
In addition to these comparisons, our proposed method demonstrates robustness to missing modalities, as evidenced by its comparison with the baseline in Section 3.6. Even in scenarios where only a single modality is available, the model’s performance is reasonably close to that of standalone unimodal models.
@@
Proposed method PlantCLEF2015 956
@@ 5 Conclusions
This study has addressed a critical task in agriculture and ecology: plant classification. By proposing a novel approach in this domain – a multimodal DL model utilizing four plant organs automatically fused via the MFAS algorithm, the study has demonstrated high performance, outperforming other state-of-the-art models despite the smaller size of the model. This underscores the effectiveness of multimodality and an optimal fusion strategy. Moreover, the model exhibits robustness to missing modalities, even when only a single modality is available. We believe that this proposed approach opens up a promising direction in plant classification research. This highlights the need for further exploration through the development of more sophisticated algorithms capable of handling larger numbers of species, thereby unlocking the full potential of this approach. A future research could also consider incorporating a multimodal fusion of vision transformers instead of CNNs for plant classification.

### v2

@@ 3.2 Finding Final Architecture
flower, leaf, fruit and stem unimodal models, respectively. $A$ – first activation, $I$ – intermediate dense layer,
$O$ – output layer, $BN$$N$-th inverted residual block.
@@
The F1macro values were achieved by the models after 2 epochs during the search process (Section 2.4.2).
@@
Figure 5: Validation F1macro scores for Sampled Architectures across Epochs
@@ 3.3 Final Model
To train the final model, we adjust various hyperparameters of the second architecture from Table 5. The full list of hyperparameters is presented in Table 6. Surprisingly, neither full randomization nor augmentations led to improved performance during multimodal model training. As a result, we continue using cached batches shuffled in buffers of 12 each epoch, as done during the MFAS execution (Section 2.4.2). Additionally, fine-tuning the model did not enhance generalization, and is therefore omitted.
Hyperparameter Value
Optimizer Adam
Initial LR $5^{-4}$
LR decay rate $0.9$
LR decay steps 200
Batch size 256
Epochs 100
ES patience 10
Classifier dropout 0.4
Fusion dropouts [0.0, 0.0, 0.0, 0.4]
Fusion neurons [512, 512, 512, 512]
MD 0.0 or 0.125
Fine-tuning False
Retraining on combined training and validation sets:
ES False
Epochs 100
Hyperparameters remain unchanged during retraining on combined
datasets unless specified otherwise.
Table 6: Hyperparameters of the Final Model
With all the aforementioned adjustments, the model reached a validation F1macro score of 68.62%. Figure 6 illustrates the increase in F1macro for this configuration following hyperparameter tuning. To prepare the model for evaluation, we merge the training and validation sets and train two versions of the model—one with MD and one without it, as discussed in Section 2.4.3.
Figure 6: Validation F1macro across Epochs before and after Final Model Hyperparameter Tuning
@@ 3.4 Comparison with the Baseline
We collect the performance metrics for both versions of the proposed model and the baseline on the test set, which are displayed in Table 7.
Proposed Model Proposed Model with MD Baseline
Accuracy 0.8261 0.7856 0.7228
Top-5 accuracy 0.9546 0.9219 0.8982
Top-10 accuracy 0.9754 0.9528 0.9420
Precisionmacro 0.7838 0.7397 0.6652
Recallmacro 0.7743 0.7130 0.6284
F1macro 0.7595 0.7024 0.6224
Table 7: Performance Metrics for Proposed Model and Baseline
These results indicate that both versions of the proposed model significantly outperform the baseline across all explored metrics, suggesting that our automated fused model is substantially more effective and reliable for plant classification on the given dataset compared to a straightforward late fusion approach. The improvement is particularly notable in the model without MD, achieving, for instance, +0.1459 in recall and +0.1371 in F1 score. Conversely, a more modest improvement is observed in the model with MD, with, for example, +0.0846 in recall and +0.08 in F1 score. This suggests that, overall, the model without MD generalizes better when evaluated over the entire test set, where modalities in instances are missing only if they are entirely absent in their corresponding classes across all splits. However, in real-world scenarios, the available set of modalities may differ from those represented in certain classes in our dataset. For this reason, we assess the models on subsets of modalities in Section 3.5.
@@ 3.4.1 McNemar’s Test Results
We also conducted McNemar’s tests to compare the performance of two model pairs: (i) the proposed model versus the baseline and (ii) the proposed model with MD versus the baseline. The contingency tables of model predictions are provided in Table 8.
The McNemar’s tests yielded statistics of $\chi^{2}_{1}=534.12$ and $\chi^{2}_{1}=228.74$, both of which are statistically significant ($p<0.001$). These results indicate a significant difference in performance metrics between each version of our proposed model and the baseline.
Without MD With MD
$n_{00}$: 1197 $n_{01}$: 281 $n_{00}$: 1468 $n_{01}$: 354
$n_{10}$: 1159 $n_{11}$: 5863 $n_{10}$: 888 $n_{11}$: 5790
@@
Table 8: Contingency Tables of Proposed Model with/without MD and Baseline
3.5 Comparison on Subsets of Modalities
@@
One can observe from Table 4 and Table 7 that both our proposed model and late fusion significantly outperform each individual unimodal model. However, these observations do not fully consider the fact that the unimodal models are evaluated on isolated unimodal datasets, whereas the multimodal model is evaluated on a multimodal dataset containing missing modalities and duplications. To address this and in order to have a fair comparison, we evaluate both our proposed models and the unimodal models on isolated modalities from the multimodal dataset. Furthermore, while cases where a single or all organs are available, represent two extremes, there exists a considerable number of image combinations within the dataset where 2–3 modalities are available. Therefore, it is essential to evaluate the performance of our proposed model and the baseline on all possible subsets of 2–3 modalities. Finally, we assess our models in cases where all modalities are available. Table 9 summarizes the results of these evaluations.
Modalities # of Predictions Proposed Model Proposed Model with MD Baseline
Flower 8097 0.3315** 0.5512** 0.5663
Leaf 6714 0.0535** 0.2133* 0.2012
Fruit 4614 0.0000** 0.1259** 0.2894
Stem 3531 0.0000** 0.0374** 0.1203
Flower, Leaf 6325 0.5120** 0.6285** 0.5159
Flower, Fruit 4357 0.2987** 0.5332 0.4574
Flower, Stem 3229 0.3384** 0.5684** 0.3655
Leaf, Fruit 4130 0.0604** 0.3665 0.3512
Leaf, Stem 3420 0.0674** 0.3588** 0.2686
Fruit, Stem 2675 0.0000** 0.2933** 0.3370
Flower, Fruit, Leaf 3887 0.5814 0.6004** 0.4787
Flower, Stem, Leaf 3129 0.5641* 0.6534** 0.4248
Flower, Fruit, Stem 2450 0.4540** 0.6114** 0.4197
Leaf, Fruit, Stem 2611 0.1289** 0.4921** 0.4092
Flower, Leaf, Fruit, Stem 2397 0.7197** 0.6768** 0.4665
* and ** denote statistical significance, with $p<0.05$ and $p<0.001$, respectively, as determined by
McNemar’s test when compared to the baseline.
@@
Table 9: Comparison of F1macro for Proposed and Baseline Models on Subsets of Modalities
It is important to note that the metrics in Table 9 account for the absence of target modalities by excluding instances that miss any of the required modalities.
The results indicate that the baseline significantly outperforms our proposed model without MD in most cases. The proposed model exhibits superior metrics only in 3 out of 4 scenarios involving 3 modalities and when all modalities are available; however, one of the cases with 3 modalities lacks statistical significance, and another shows lower significance compared to other comparisons. Notably, there are instances where our model achieves a 0% F1 score, specifically in cases involving only fruits and/or stems, while subsets that include leaves demonstrate higher scores, and those incorporating flowers yield significantly better results than any other subsets. The distribution of scores among the unimodal models suggests a different trend, with the informal notation $F1_{flower}>F1_{fruit}>F1_{leaf}>F1_{stem}$ in unimodal models versus $F1_{flower}>F1_{leaf}>F1_{stem}\geq F1_{fruit}$ in the proposed model, implying that the proposed model may be sensitive to the imbalance in modalities, leaning more towards learning from the more common modalities. Consequently, to enhance performance on subsets with less represented modalities, techniques such as augmenting the data with additional instances of underrepresented modalities or applying weighted loss based on the modalities of a predicted instance should be considered. Overall, the findings suggest that the proposed model without MD is more effective when the majority of modalities are available; however, since it consistently trains with the maximum number of available modalities, this approach ultimately leads to reduced robustness when modalities are missing.
Conversely, the proposed model with MD significantly outperforms the baseline in the majority of cases, indicating the effectiveness of the MD technique in enhancing robustness to missing modalities. Surprisingly, its performance with a single flower modality is very similar to that of the baseline and even better with the leaf modality, despite the baselines, in this cases, being standalone unimodal models specifically designed for individual modalities. This observation highlights the model’s ability to resist noise from idle modality-specific parts. Notably, this version of our model surpasses the one without MD in all cases except for the scenario where all modalities are available. This suggests that the robustness to missing modalities afforded by MD comes at a slight cost to performance across all modalities; the model without MD consistently trains on all available modalities simultaneously, facilitating the learning of interconnections between them, while the model with MD rarely encounters the full set of modalities during training.
Overall, our approach outperforms the late fusion of unimodal models. However, if robustness to missing modalities is crucial, it is essential to employ MD or similar techniques. Conversely, to maximize performance when the full set of available modalities is present, the model should consistently train with all the available modalities visible.
@@ 4 Discussion
Our proposed multimodal DL model, which utilizes plant organ images fused through the MFAS algorithm, demonstrates high effectiveness in automating plant classification, achieving an accuracy of 82.61% and significantly surpassing the defined baseline. While a strict comparison with previous studies is not feasible due to unique aspects of our data preprocessing and methodology, Table 10 indicates that our model achieves the highest accuracy among similar research. To be more specific, for example, Ghazi et al. (2017) employed all organs from the PlantCLEF2015 dataset. Despite utilizing larger and more sophisticated classifiers, along with advanced scoring methods, they achieved an accuracy of 80.18%. Despite our architecture being relatively small (10 million parameters), the incorporation of multimodal fusion allowed us to outperform their model, highlighting the strength of multimodality and the advantages of an optimized fusion strategy of our proposed method. Similarly, de Lutio et al. (2021) sampled a large dataset comprising 56608 images of 977 species from the iNaturalist database. Utilizing images and spatio-temporal context, their multimodal model achieved accuracies of 79.12% without satellite imagery and 79.73% with it. While a direct comparison is challenging due to the use of different datasets, our approach, incorporating various organ modalities, yielded higher results. Furthermore, Lee et al. (2018), also using the PlantCLEF2015 dataset, achieved an accuracy of 68.5%. Ge et al. (2016) and Nguyen et al. (2016) extracted flower species from the PlantCLEF2015 dataset and achieved an accuracy of 52.1% and 67.45% respectively. All these results are significantly lower than ours, underscoring the effectiveness of our approach.
In addition to these comparisons, our proposed method demonstrates robustness to missing modalities, as evidenced by its comparison with the baseline in Section 3.5.
@@
Proposed method PlantCLEF2015 979
@@ 5 Conclusions
This study has addressed a critical task in agriculture and ecology: plant classification. By proposing a novel approach in this domain—a multimodal DL model utilizing four plant organs automatically fused via the MFAS algorithm, the study has demonstrated high performance, outperforming other state-of-the-art models despite the smaller size of the model. This underscores the effectiveness of multimodality and an optimal fusion strategy. Moreover, with the MD technique, the model exhibits robustness to missing modalities, even when only a single modality is available. We believe that this proposed approach opens up a promising direction in plant classification research. This highlights the need for further exploration through the development of more sophisticated algorithms capable of handling larger numbers of species, thereby unlocking the full potential of this approach. A future research could also consider incorporating a multimodal fusion of vision transformers instead of CNNs for plant classification.